NvidiaIntely (nvidiaintely):

A sample of Strontium-90 is found to have decayed to one-eighth of its original amount after 87.3 years. What is the half-life of strontium-90? 87.3 years 43.7 years 29.1 years 10.9 years

OpenStudy (kevin):

is it 1/8 ?

NvidiaIntely (nvidiaintely):

What you mean?

OpenStudy (kevin):

about one-eighth

NvidiaIntely (nvidiaintely):

But thats not in the choices. 87.3 years 43.7 years 29.1 years 10.9 years

OpenStudy (kevin):

I know, one-eighth is same 1/8 ?

NvidiaIntely (nvidiaintely):

yes.

OpenStudy (kevin):

Nt = 1/8 No Nt/No = 1/8 Nt/No = 1/2^(t/delta t) 1/8 = 1/2^(87.3/delta t) delta t = 29.1

NvidiaIntely (nvidiaintely):

Oh okay, thanks.

OpenStudy (kevin):

:)